OpenStudy (usarmy3947):

What is the value of the expression 7 + 52 × (7 – 3)?

OpenStudy (usarmy3947):



OpenStudy (usarmy3947):

74 107 128 179

OpenStudy (anonymous):

just solve it 7+52x(7-3) OOO,7-3 is 4. 7-52x4 7-25x4 7-100 -93

OpenStudy (usarmy3947):

thats not a answer choice

OpenStudy (anonymous):

oh, 7PLUS100, the answer is 107

OpenStudy (usarmy3947):

thanks for your help @BeckieBird

OpenStudy (ja1):

I thought that was an exponent *scratches head*

OpenStudy (anonymous):

it is an exponent, do order of operations

OpenStudy (usarmy3947):

@JA1 can u help with this one

OpenStudy (usarmy3947):

What number is missing in the expression

OpenStudy (usarmy3947):



OpenStudy (usarmy3947):

i know i have asked this one but really need help with it

OpenStudy (ja1):

Okay well as for the "acceptable" method I got nothing, I can solve this but I don't know the correct method that is taught, is that ok?

OpenStudy (usarmy3947):

of course if u want just make a guess

OpenStudy (ja1):

No not a guess, I just don't know the right steps, my way of doing it is as follows:

OpenStudy (ja1):

Wait to verify it's: \[3^2x \left(?+3\right)=63\]

OpenStudy (usarmy3947):

please hurry my tutor wants me to be done in 1 min. and give the results to my parents and if that question isn't answered my parents will whoop me

OpenStudy (ja1):

sorry for the delay by browser is glitchy today, anyways I need to know if that is the question?

OpenStudy (usarmy3947):

yes it is

OpenStudy (ja1):

okay well first I do the steps except for parenthesis, so I get: 9*3=63 which is wrong so I need to figure what number times 9 is 63, I know it's 7 off the top of my head, and thus the number is 4 since 4+3 = 7!

OpenStudy (ja1):

So you get: \[3^2(7 + 3) = 63\]

OpenStudy (usarmy3947):

thanks so much @JA1 ur a life saver

OpenStudy (ja1):

I thought "x" was a variable xD now I get that it's a multiplication sign *smacks head*

OpenStudy (ja1):

No problem, btw try using "*" from now on to show multiplication ;) it gets confused for algebra

OpenStudy (usarmy3947):

lol thanks for the heads up :)

OpenStudy (ja1):

No prob, good luck!

OpenStudy (usarmy3947):

i got 14 out of 20

OpenStudy (ja1):

That's a 61 D... might want to study a bit harder next time :)

OpenStudy (usarmy3947):

ah my dad is coming with the belt now

OpenStudy (ja1):

*Weeps for Army3947*

OpenStudy (usarmy3947):

im scared

OpenStudy (ja1):

Plead guilty and promise to study harder ;) it "should" work

OpenStudy (usarmy3947):

no it wont he doesn't have mercy he's a crazy man he's like a helicopter with that belt

OpenStudy (anonymous):

215. Just use a calculator. LOL

OpenStudy (anonymous):

Just follow the order of operations. First evaluate exponents, then numbers within parentheses, then multiplication and division, and then addition and subtraction. In this case, you must solve 7 - 3, because it is within parentheses, and then multiply that answer by 52, since multiplication comes next. Because addition comes last, you must add that value to 7 and you're done! If you want to solve this kind of problem on your own in the future, just remember the order of operations and you'll be fine. :)